### v1

@@ 3 Method
Figure 1 shows an overview of the pipeline for measuring the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ of cool stars in the LAMOST LRS DR10. Four swimlanes in the figure represent the four modules of the pipeline, which are responsible for building the target and reference dataset, deriving the ISM residual spectra of the target dataset, measuring the DIBs in the residual spectra, and quality control of the final measurements. The details of each module are described in the following subsections.
Figure 1: Schematic flowchart of the pipeline for measuring the DIBs at 5780, 5797 and 6614 Å of cool stars in the LAMOST LRS DR10. Four colors encode the four parts of the pipeline, and the output of each modular function is used for the input of the next one.
@@ 3.1 Building the Dataset
For the cool stars, the regions nearby DIBs are usually blended with the stellar absorption lines, which need to be thoroughly removed as much as possible. Inspired by the cases for the success in detecting the DIBs in the spectra of cool stars (Kos et al. 2013; Yuan & Liu 2012; Lan et al. 2015; Vogrinčič et al. 2023), we build two datasets, i.e., the target dataset in which the DIBs are likely to be embedded in the spectra, and the reference dataset in which the spectra are expected to contain the pure stellar features within the region of the DIB to be measured.
At first, both the target and reference datasets are filtered by the basic data quality control for the better stellar parameter matching in Section 3.2. They are required to have valid stellar parameters within the error of 100 K, 0.2 dex, 0.1 dex and 10 $\rm km\ s^{-1}$ for $T_{\rm eff}$, $\log g$, [Fe/H] and radial velocity (RV), respectively.
In general, given that the absorption depths of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are commonly larger than one percent of continuum (Vogrinčič et al. 2023), we conservatively select the spectra with the signal-to-noise ratio (SNR) in $r$ band (the keyword snr_r or SNR_r, which is derived from the average SNR of the pixels within the $r$ band, i.e., from 5600 to 6800 Å) greater than 100 as our target dataset. In that the reference dataset are used to subtract the components that do not contribute to the DIB in the target dataset, they are supposed to meet the following conditions to ensure that the interstellar clouds in the foreground are as thin as possible: (i) snr_r $>50$; (ii) $\rm E(B-V)<0.03\ mag$; (iii) Galactic latitude, $|b|>60^{\circ}$. There are 2,188,240 spectra in the target dataset and 424,164 spectra in the reference dataset after the above filtering. The distribution of three atmospheric parameters for the two datasets are shown in Figure 2. The JS divergences (Menéndez et al. 1997), which range from 0 to 1 and are often used to measure the similarity between two distributions, of $T_{\rm eff}$, $\log g$ and [Fe/H] between the target and reference dataset are 0.0589, 0.0410 and 0.0278, respectively. These values closer to zero indicate it is of high similarity between the two datasets, which makes the reference dataset matchable for the subtraction of the stellar features in the target dataset.
Figure 2: Distribution of the three atmospheric parameters, i.e., $T_{\rm eff}$ (left panel), $\log g$ (middle panel) and [Fe/H] (right panel), for the target and reference datasets. The JS divergences of the three parameters between the two datasets are annotated in each panel.
@@ 3.2 Getting the ISM Residual Spectra
An ISM residual spectrum is the integral signals of DIB at the line-of-sight of a target object, so that an observed spectrum in the target dataset can be viewed as the product of the ISM residual and the stellar components. Although it is impossible to resolve the authentic stellar features out of a target spectrum, one of the workarounds is stacking the neighbor spectra of the target in the reference dataset to approximate the stellar ingredients we call the template spectrum (Kos et al. 2013). Then, the ISM residual spectrum can be obtained by the ratio of the target spectrum to the template spectrum from the stacked closest spectra, as can be seen in Figure 3.
Figure 3: An example of measuring the DIBs $\lambda 5780$ and $\lambda 5797$ for a target spectrum. From top to bottom, (i) the first panel shows the target spectrum to be measured, with the red texts indicating its obsid (the unique ID of the spectrum in the LAMOST catalog) and atmospheric parameters. (ii) The second panel provides the first 30 best neighbor spectra of the target spectrum, which are used to build the template spectrum displayed in the third panel. The purple-shaded areas in the first and second panels are the masked regions that do not participate in the similarity computation. (iii) The third panel illustrates the target and template spectrum after local normalization. The most prominent stellar absorptions which may affect the DIB extraction are annotated in the panel. (iv) The fourth panel presents the ISM residual spectrum (in black) of the target and the fit (in red) of the DIBs $\lambda 5780$ and $\lambda 5797$ by MCMC. The shallow DIBs $\lambda 5762$ and $\lambda 5815$ in the ISM residual spectrum are also annotated but not fitted in the panel.
@@ 3.2.1 Matching Neighbors
The quality of the best neighbor matching determines whether the stellar components in the target spectrum can be cleanly deducted. It is essential to ensure that these neighbors are as close to the target possible. We impose constraints on the selection of the best neighbor samples based on the atmospheric parameter space, match number, SNR, and similarity in spectral morphology.
We make a preliminary selection within a certain range of atmospheric parameters. For the samples in the reference dataset that satisfy $\Delta T_{\rm eff}<100$ K, $\Delta\log g<0.5$ dex, $\rm\Delta[Fe/H]<0.2$ dex with respect to the target, if the number of those samples is no more than 100, we drop them out. Then, to balance the computation efficiency and match accuracy, the top 1000 samples (if enough) with the highest SNR_r as the neighbor candidates of the target are retained.
However, the proximity of atmospheric parameter space cannot guarantee the similarity of spectral morphology due to the measurement error of parameters. We further adopt the reciprocal of Euclidean distance between the target and neighbor to represent the similarity in spectral morphology. Each continuum-normalized spectrum is processed for the alignment of similarity computation as follows: (i) mask the bad flux pixels according to the ormask flag; (ii) shift the spectrum to the rest frame; (iii) convert the wavelength from vacuum to air; (iv) truncate the spectrum to [4500, 7500] Å to avoid the bias of Balmer series at the blue end and the low response efficiency at the red end, and then rebin the spectrum with a step of 1 Å; (v) mask the regions that do not contribute to the metal lines, including H $\alpha$ , H $\beta$ , strong DIBs $\lambda 5780$, $\lambda 5797$, $\lambda 6283$ and $\lambda 6614$ (see Table 1 for their detailed wavelength ranges). Ultimately, the top 25% similar candidate samples, i.e., those with no fewer than 25 and at most 250, are selected as the closest neighbors prepared for the stacked template spectrum in Section 3.2.3. For instance, the first 30 closest neighbors of a target spectrum of obsid = 837603202 can be seen in the second panel of Figure 3.
Table 1: The masked region used to match neighbor spectra in Section 3.2.1
@@ 3.2.2 Preprocessing
Similar to Section 3.2.1, a series of preprocessing steps are required to align the target and its neighbors before extracting the ISM residual spectrum of the target. These preprocessing steps still include removing the bad flux pixels, shifting the spectrum to the rest frame and converting the wavelength from vacuum to air. Unlike Section 3.2.1, the re-normalized flux within the local region instead of the continuum-normalized flux is utilized. Although the continuum-normalized flux provided by LASP meets the requirement of spectral morphology matching in Section 3.2.1, it is too coarse to measure the weak DIBs when anchoring the local region of the DIBs. Therefore, a more careful local normalization for the flux within the local DIB region needs to be performed.
Initially, the wavelength range of the local DIB region is determined. For the DIB $\lambda 5780$ and $\lambda 5797$, they are merged into a single local region, i.e., [5755, 5855] Å  owing to their close central wavelengths. For the DIB $\lambda 6614$, the blue end keeps away from the H $\alpha$ line to alleviate the influence of the quite strong H $\alpha$ feature on the blue wing of the DIB $\lambda 6614$. In addition, the red end of the local region is moderately extended to compensate for the accuracy of the local continuum normalization. The final selection for the DIB $\lambda 6614$ is [6590, 6690] Å. Subsequently, the flux in these two local regions is linearly interpolated and rebinned to an interval of 0.8 Å, resulting in a total of 125 pixels. We refer to the method by Zhao et al. (2021a) as the local normalization of the flux within the selected local DIB region. The only difference is that we use a fifth-order polynomial curve instead of a second-order one to fit the local spectrum because of the cool stars in this work. A 20-iteration continuum fitting is performed as follows: (i) fit the local spectrum by a fifth-order polynomial curve; (ii) compute the differences of the local spectrum to the fitting curve, as well as their standard deviation; (iii) replace the flux values with the fitted polynomial ones if the differences between them are larger than five times of the standard deviation when the pixels are located above the polynomial curve, or 0.5 times when the pixels are below the polynomial curve; (iv) use the remaining and replaced pixels as the new local spectrum to repeat the above steps. The final normalized spectrum is obtained by dividing the last local spectrum by the last fitted continuum. Figure 4 provides two examples of the local renormalization for the target spectra with obsid = 894515122 and obsid = 794616199. It is clear that the local renormalization can effectively correct the continuum level of some absorption features.
@@
The flux error is estimated while performing the local normalization because the flux error after local normalization is not only a key input parameter for the subsequent MCMC fitting in Section 3.3, but also an important indicator for evaluating the quality of the ISM residual spectra in Section 4. Although the LAMOST provides the error of the flux in a format of the inverse variance (1/error2), many works, e.g., Xiang et al. (2015) and Ho et al. (2017) have reported this error is overestimated. Ho et al. (2017) and Zhang et al. (2020) have given the same approximations of the inverse variance of the flux after normalization, which can be expressed as Equation 1,
@@
where $\rm F_{cont}$ is the fitted continuum and $\rm ivar_{rel}$ is the inverse variance of the flux. The final re-normalized flux error equals the inverse of the square root of $\rm ivar_{norm}$.
@@ 3.2.3 Isolating the ISM residual spectra
Once the locally normalized spectra of the target and its best neighbors are ready, the high-quality template spectrum for the target can be obtained by averaging those spectra of the best neighbors weighted by their SNR_r, which characterizes the stellar features in the target spectrum. Then, the ISM residual spectrum of the target is isolated by dividing the target spectrum by the template spectrum, as shown in the third and fourth panels of Figure 3. By comparing the template spectrum to the spectral standard star Arcturus (a K1.5 III star with the luminosity of 170 $L_{\odot}$) using an interactive database (Lobel 2011) called SpectroWeb (http://spectra.freeshell.org/spectroweb.html), we also list the most prominent stellar lines that may contaminate the DIBs in the template spectrum in the third panel of Figure 3.
The error of the ISM spectrum is deduced by the error propagation as Equation 2,
@@
3.3 Measuring the DIBs
Most of the DIB profiles are fitted by the symmetric Gaussian function, e.g., Lan et al. (2015) and Zhao et al. (2021a), but some studies based on the early-type stars in high resolution have found that the profiles of some DIBs are asymmetric and multiple components. For instance, Galazutdinov et al. (2008) reported the DIB $\lambda 6614$ displays a triple-peak fine structure. Considering the low resolution of LAMOST, the instrumental broadening effect (see Appendix A) dominates the shape of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ and their fine structures are blurred. In such a resolution, Yuan & Liu (2012), Lan et al. (2015) and Farhang et al. (2019) have proven the Gaussian function is still a good approximation for the profile of those DIBs. Thus, we adopt the Gaussian function to fit the DIBs in the ISM residual spectra. The process of measuring the DIBs consists of three steps: (i) pre-detect the central wavelength and depth of the DIB to determine whether there is a DIB signal; (ii) use the parameters obtained in the first step and the fixed Gaussian widths (3.85, 2 and 2 Å for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively) as the initial values of the least-square curve fitting to get the three parameters of the Gaussian profile, i.e., the central wavelength, depth and Gaussian width; (iii) further optimize the parameters obtained in the second step using the MCMC method to get the final fitting results. An MCMC fitting demo of the DIBs $\lambda 5780$ and $\lambda 5797$ is presented in the fourth panel of Figure 3.
3.3.1 Pre-detecting
In the context of big data mining, pre-detection can not only quickly identify the samples probably contain DIBs, but also provide reasonable initial values for the curve fitting in Section 3.3.2. We take the rest-frame central wavelength and FWHM of the DIBs reported by Fan et al. (2019) into account, and set the wavelength range of the pre-detection segment to be about ±$\pm 300\ \rm km\ s^{-1}$ around the central wavelength of the DIB. For the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, their pre-detection segments are [5775, 5785], [5793, 5801] and [6608, 6620] Å, respectively. Such a wide wavelength range is chosen to deal with those potential DIBs with large RVs or broad profiles. If the maximum depth of the ISM residual spectrum within the pre-detection segment exceeds 1/SNR_r of the target spectrum, where SNR_r is the average SNR within the $r$ band that covers the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, we proceed to the next step of curve fitting. The maximum depth and the wavelength corresponding to it are taken as the initial depth and central wavelength of the DIB profile. Otherwise, we drop out the target.

### v2

@@ 3 Method
Figure 1 shows an overview of the pipeline for measuring the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ of cool stars in the LAMOST LRS DR10. Four swimlanes in the figure represent the four modules of the pipeline, which are responsible for building the target and reference dataset, deriving the ISM residual spectra of the target dataset, measuring the DIBs in the residual spectra, and quality control of the final measurements. The details of each module are described in the following subsections.
Figure 1: Schematic flowchart of the pipeline for measuring the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ of cool stars in the LAMOST LRS DR10. Four colors encode the four parts of the pipeline, and the output of each modular function is used for the input of the subsequent one.
@@ 3.1 Building the dataset
For the cool stars, the regions nearby DIBs are usually blended with the stellar absorption lines, which need to be thoroughly removed as much as possible. Inspired by the cases for the success in detecting the DIBs in the spectra of cool stars (Kos et al. 2013; Yuan & Liu 2012; Lan et al. 2015; Vogrinčič et al. 2023), we build two datasets, that is, the target dataset in which the DIBs are likely to be embedded in the spectra, and the reference dataset in which the spectra are expected to contain the pure stellar features within the region of the DIB to be measured.
At first, both the target and reference datasets are filtered by the basic data quality control for the better stellar parameter matching in Sect. 3.2. They are required to have valid stellar parameters within the error of 100 K, 0.2 dex, 0.1 dex, and 10 $\rm km\ s^{-1}$ for $T_{\rm eff}$, $\log g$, [Fe/H], and radial velocity (RV), respectively.
In general, given that the absorption depths of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are commonly larger than one percent of continuum (Vogrinčič et al. 2023), we conservatively select the spectra with the signal-to-noise ratio in $r$ band (the keyword snrr in the LAMOST catalog, hereafter S/N, which is derived from the average signal-to-noise ratio of the pixels within the $r$ band, i.e., from 5600 to 6800 Å) greater than 100 as our target dataset. In that the reference dataset are used to subtract the components that do not contribute to the DIB in the target dataset, they are supposed to meet the following conditions to ensure that the interstellar clouds in the foreground are as thin as possible: (i) S/N $>50$; (ii) $\rm E(B-V)<0.03\ mag$; (iii) Galactic latitude, $|b|>60^{\circ}$. There are 2,188,240 spectra in the target dataset and 424,164 spectra in the reference dataset after the above filtering. The distribution of three atmospheric parameters for the two datasets are shown in Fig. 2. The JS divergences (Menéndez et al. 1997), which range from 0 to 1 and are often used to measure the similarity between two distributions, of $T_{\rm eff}$, $\log g$, and [Fe/H] between the target and reference dataset are 0.0589, 0.0410, and 0.0278, respectively. These values closer to zero indicate it is of high similarity between the two datasets, which makes the reference dataset matchable for the subtraction of the stellar features in the target dataset.
Figure 2: Distribution of the three atmospheric parameters, namely, $T_{\rm eff}$ (left panel), $\log g$ (middle panel), and [Fe/H] (right panel), for the target and reference datasets. The JS divergences of the three parameters between the two datasets are annotated in each panel.
@@ 3.2 Getting the ISM residual spectra
An ISM residual spectrum is the integral signals of DIB at the line-of-sight of a target object, so that an observed spectrum in the target dataset can be viewed as the product of the ISM residual and the stellar components. Although it is impossible to resolve the authentic stellar features out of a target spectrum, one of the workarounds is stacking the neighboring spectra of the target in the reference dataset to approximate the stellar ingredients we call the template spectrum (Kos et al. 2013). Then, the ISM residual spectrum can be obtained by the ratio of the target spectrum to the template spectrum from the stacked closest spectra, as can be seen in Fig. 3.
Figure 3: Example measurement of the DIBs $\lambda 5780$ and $\lambda 5797$ from a target spectrum. From top to bottom, (i) the first panel shows the target spectrum to be measured, with the red text indicating its obsid (the unique ID of the spectrum in the LAMOST catalog) and atmospheric parameters. (ii) The second panel provides the first 30 best neighboring spectra of the target spectrum, which are used to build the template spectrum displayed in the third panel. The purple-shaded areas in the first and second panels are the masked regions that do not participate in the similarity computation. (iii) The third panel illustrates the target and template spectrum after local normalization. The most prominent stellar absorptions that may affect the DIB extraction are annotated in the panel. (iv) The fourth panel presents the ISM residual spectrum (in black) of the target and the fit (in red) of the DIBs $\lambda 5780$ and $\lambda 5797$ by MCMC. The shallow DIBs $\lambda 5762$ and $\lambda 5815$ in the ISM residual spectrum are also annotated but not fitted in the panel.
@@ 3.2.1 Matching neighbors
The quality of the best neighbor matching determines whether the stellar components in the target spectrum can be cleanly deducted. It is essential to ensure that these neighbors are as close to the target possible. We impose constraints on the selection of the best neighboring samples based on the atmospheric parameter space, match number, S/N, and similarity in spectral morphology.
We make a preliminary selection within a certain range of atmospheric parameters. For the samples in the reference dataset that satisfy $\Delta T_{\rm eff}<100$ K, $\Delta\log g<0.5$ dex, and $\rm\Delta[Fe/H]<0.2$ dex with respect to the target, if the number of those samples is no more than 100, we drop them out. Then, to balance the computation efficiency and match accuracy, the top 1000 samples (if enough) with the highest S/N in $r$ band as the neighboring candidates of the target are retained.
However, the proximity of atmospheric parameter space cannot guarantee the similarity of spectral morphology due to the measurement error of parameters. We further adopt the reciprocal of Euclidean distance between the target and the neighbor to represent the similarity in spectral morphology. Each continuum-normalized spectrum is processed for the alignment of similarity computation as follows: (i) mask the bad flux pixels according to the ormask flag; (ii) shift the spectrum to the rest frame; (iii) convert the wavelength from vacuum to air; (iv) truncate the spectrum to [4500, 7500] Å to avoid the bias of Balmer series at the blue end and the low response efficiency at the red end, and then rebin the spectrum with a step of 1 Å; (v) mask the regions that do not contribute to the metal lines, including H $\alpha$ , H $\beta$ , strong DIBs $\lambda 5780$, $\lambda 5797$, $\lambda 6283$, and $\lambda 6614$ (see Table 1 for their detailed wavelength ranges). Ultimately, the top 25% similar candidate samples, namely those with no fewer than 25 and at most 250, are selected as the closest neighbors prepared for the stacked template spectrum in Sect. 3.2.3. For instance, the first 30 closest neighbors of a target spectrum of obsid = 837603202 can be seen in the second panel of Fig. 3.
Table 1: Masked region used to match neighboring spectra in Sect. 3.2.1
@@ 3.2.2 Preprocessing
Similar to Sect. 3.2.1, a series of preprocessing steps are required to align the target and its neighbors before extracting the ISM residual spectrum of the target. These preprocessing steps still include removing the bad flux pixels, shifting the spectrum to the rest frame, and converting the wavelength from vacuum to air. Unlike Sect. 3.2.1, the renormalized flux within the local region instead of the continuum-normalized flux is utilized. Although the continuum-normalized flux provided by LASP meets the requirement of spectral morphology matching in Sect. 3.2.1, it is too coarse to measure the weak DIBs when anchoring the local region of the DIBs. Therefore, a more careful local normalization for the flux within the local DIB region needs to be performed.
Initially, the wavelength range of the local DIB region is determined. For the DIB $\lambda 5780$ and $\lambda 5797$, they are merged into a single local region, namely [5755, 5855] Å  owing to their close central wavelengths. For the DIB $\lambda 6614$, the blue end keeps away from the H $\alpha$ line to alleviate the influence of the quite strong H $\alpha$ feature on the blue wing of the DIB $\lambda 6614$. In addition, the red end of the local region is moderately extended to compensate for the accuracy of the local continuum normalization. The final selection for the DIB $\lambda 6614$ is [6590, 6690] Å. Subsequently, the flux in these two local regions is linearly interpolated and rebinned to an interval of 0.8 Å, resulting in a total of 125 pixels. We refer to the method by Zhao et al. (2021a) as the local normalization of the flux within the selected local DIB region. The only difference is that we use a fifth-order polynomial curve instead of a second-order one to fit the local spectrum because of the cool stars in this work. A 20-iteration continuum fitting is performed as follows: (i) fit the local spectrum by a fifth-order polynomial curve; (ii) compute the differences of the local spectrum to the fitting curve, as well as their standard deviation; (iii) replace the flux values with the fitted polynomial ones if the differences between them are larger than five times of the standard deviation when the pixels are located above the polynomial curve, or 0.5 times when the pixels are below the polynomial curve; (iv) use the remaining and replaced pixels as the new local spectrum to repeat the above steps. The final normalized spectrum is obtained by dividing the last local spectrum by the last fitted continuum. Figure 4 provides two examples of the local renormalization for the target spectra with obsid = 894515122 and obsid = 794616199. It is clear that the local renormalization can effectively correct the continuum level of some absorption features.
@@
The flux error is estimated while performing the local normalization because the flux error after local normalization is not only a key input parameter for the subsequent MCMC fitting in Sect. 3.3, but also an important indicator for evaluating the quality of the ISM residual spectra in Sect. 4. Although the LAMOST provides the error of the flux in a format of the inverse variance (1/error2), many works, such as Xiang et al. (2015) and Ho et al. (2017) have reported this error is overestimated. Ho et al. (2017) and Zhang et al. (2020) have given the same approximations of the inverse variance of the flux after normalization, which can be expressed as Eq. (1),
@@
where $\rm F_{cont}$ is the fitted continuum and $\rm ivar_{rel}$ is the inverse variance of the flux. The final renormalized flux error equals the inverse of the square root of $\rm ivar_{norm}$.
@@ 3.2.3 Isolating the ISM residual spectra
Once the locally normalized spectra of the target and its best neighbors are ready, the high-quality template spectrum for the target can be obtained by averaging those spectra of the best neighbors weighted by their S/N in $r$ band, which characterizes the stellar features in the target spectrum. Then, the ISM residual spectrum of the target is isolated by dividing the target spectrum by the template spectrum, as shown in the third and fourth panels of Fig. 3. By comparing the template spectrum to the spectral standard star Arcturus (a K1.5 III star with the luminosity of 170 $L_{\odot}$) using an interactive database (Lobel 2011) called SpectroWeb (http://spectra.freeshell.org/spectroweb.html), we also list the most prominent stellar lines that may contaminate the DIBs in the template spectrum in the third panel of Fig. 3.
The error of the ISM spectrum is deduced by the error propagation as Eq. (2),
@@
3.3 Measuring diffuse interstellar bands
Most of the DIB profiles are fitted by the symmetric Gaussian function, such as Lan et al. (2015) and Zhao et al. (2021a), but some studies based on the early-type stars in high resolution have found that the profiles of some DIBs are asymmetric and multiple components. For instance, Galazutdinov et al. (2008) reported the DIB $\lambda 6614$ displays a triple-peak fine structure. Considering the low resolution of LAMOST, the instrumental broadening effect (see Appendix A) dominates the shape of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ and their fine structures are blurred. In such a resolution, Yuan & Liu (2012), Lan et al. (2015), and Farhang et al. (2019) have proven the Gaussian function is still a good approximation for the profile of those DIBs. Thus, we adopt the Gaussian function to fit the DIBs in the ISM residual spectra. The process of measuring the DIBs consists of three steps: (i) pre-detect the central wavelength and depth of the DIB to determine whether there is a DIB signal; (ii) use the parameters obtained in the first step and the fixed Gaussian widths (3.85, 2, and 2 Å for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively) as the initial values of the least-square curve fitting to get the three parameters of the Gaussian profile, namely, the central wavelength, depth, and Gaussian width; (iii) further optimize the parameters obtained in the second step using the MCMC method to get the final fitting results. An MCMC fitting demo of the DIBs $\lambda 5780$ and $\lambda 5797$ is presented in the fourth panel of Fig. 3.
3.3.1 Pre-detection
In the context of big data mining, pre-detection not only allows the rapid identification of samples that probably contain DIBs, but also provides reasonable initial values for the curve fitting in Sect. 3.3.2. We take the rest-frame central wavelength and FWHM of the DIBs reported by Fan et al. (2019) into account, and set the wavelength range of the pre-detection segment to be about ±$\pm 300\ \rm km\ s^{-1}$ around the central wavelength of the DIB. For the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, their pre-detection segments are [5775, 5785], [5793, 5801], and [6608, 6620] Å, respectively. Such a wide wavelength range is chosen to deal with those potential DIBs with large RVs or broad profiles. If the maximum depth of the ISM residual spectrum within the pre-detection segment exceeds $\rm\frac{1}{S/N}$ of the target spectrum, where S/N is the average signal-to-noise ratio within the $r$ band that covers the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, we proceed to the next step of curve fitting. The maximum depth and the wavelength corresponding to it are taken as the initial depth and central wavelength of the DIB profile. Otherwise, we drop out the target.